three years ago mary was three times as old as her sister. now mary is twice as old as her sister. how old is mary now.
please show work

Answer by ptaylor(2198)

Let x=Mary's age now
And let y=her sister's age now
Then x-3=Mary's age three years ago
Now we are told that x-3=3(y-3)------------------eq1
and that:
x=2y-------------------------------------------------eq2
substitute x=2y into eq1 and we get:
2y-3=3(y-3) get rid of parens
2y-3=3y-9 add three and subtract 3y from both sides
2y-3y-3+3=3y-3y-9+3 collect like terms
-y=-6 divide both sides by -1
y=6--------------------Mary's sister's age now
substitute into eq2
x=2*6
x=12-----------------Mary's age now age now
CK
x-3=3(y-3)
12-3=3(6-3)
12-3=3*3
9=9
also
12=2*6
12=12

Hope this helps---ptaylor